OpenStudy (anonymous):

(x^2-1/4x-1/8)/(x+1/4) evaluate

OpenStudy (pfenn1):

Is this the equation?\[\left( x^2 -\frac1 4 x- \frac 18 \right)\over x+\frac 1 4\]

OpenStudy (anonymous):

yes

OpenStudy (pfenn1):

I would find the LCD of the top first\[ x^2 -\frac1 4 x- \frac 18\]what would be the LCD?

OpenStudy (pfenn1):

It would be 8. so the expression would be\[x^2 -\frac1 4 x- \frac 18=\left( 8x^2-2x-1\over 8 \right)\]

OpenStudy (anonymous):

that would be 8

OpenStudy (pfenn1):

Do the same for the expression\[x +\frac 1 4=\left( 4x+1 \over 4 \right)\]

OpenStudy (pfenn1):

Now you can write the original equation\[\left[ \left( x^2 -\frac1 4 x- \frac 18 \right)\over x+\frac 1 4 \right]=\left( 8x^2-2x-1 \over 8 \right)\left( 4 \over 4x+1 \right)=\]

OpenStudy (pfenn1):

Now factor \[8x^2-2x-1\] Hint: one of the factors will be 4x+1

OpenStudy (anonymous):

ty..trying to work it out now.....hold on one sec, please

OpenStudy (anonymous):

i got 2x-1

OpenStudy (pfenn1):

Okay, now put it all together\[\left( 8x^2-2x-1 \over 8 \right)\left( 4 \over 4x+1 \right)=\left( (2x-1)(4x+1) \over8 \right) \left( 4 \over 4x+1 \right)\] and simplify

hero (hero):

Don't know if it's right, but medal for effort.

OpenStudy (pfenn1):

What did you get for the final simplified expression?

OpenStudy (anonymous):

i keep gettt...ugh

OpenStudy (anonymous):

i keep getting something tottally different

OpenStudy (pfenn1):

Well, if I did okay up to the point above, then the final expression should be \[\frac 1 2 (2x-1)=x-\frac 1 2\]

OpenStudy (anonymous):

how did you get that???

OpenStudy (pfenn1):

From above\[\left( (2x-1)(4x+1) \over8 \right) \left( 4 \over 4x+1 \right)\]The 4x+1 term cancels and 4/8=1/2 so then\[=\frac 1 2 (2x-1)=x-\frac 1 2\]

OpenStudy (pfenn1):

Maybe someone else will check what I did and see if I made any mistakes....

OpenStudy (anonymous):

i'm checking it by doing the step i missed

OpenStudy (anonymous):

i got confused with the sign after you multiplied the recipricol of 4x+1/4

OpenStudy (anonymous):

it should be 4/4x-1

OpenStudy (anonymous):

thanks for staying with me until understood it!!!

OpenStudy (pfenn1):

I'm not sure which part you are saying should be 4/4x-1?